OpenStudy (anonymous):

Prove product of 2 odd integers is odd

OpenStudy (anonymous):

product is multiplication

OpenStudy (anonymous):

is using the fact that an even number times odd number is even is fine for the proof ?

OpenStudy (anonymous):

Let x and y be particular but arbitrarily chosen odd numbers. Then, x = 2k+1 and y = 2l + 1, for some integers k and l. We have x · y = (2k + 1) · (2l + 1) = 4kl + 2(k + l) + 1 = 2(2kl + k + l) + 1 Let p = 2kl + k + l. Since k and l are integers, p is an integer and x · y = 2p + 1 is odd.

OpenStudy (anonymous):

if so : let w,v be odd integers so using integers p,t we might write: w = 2p + 1 , v = 2t + 1 so w * v = (2p+1)(2t+1) = 4pt + 2p + 2t + 1 = 2(2pt+p+t) + 1 now if we let again x = 2pt + p +t is an integer we get that w*v is odd :)